OpenStudy (123456789):

PLz help : Prove cos(2x)^2 = cos(4x).

OpenStudy (anonymous):

let \[x=\frac{\pi}{6}\] then \[\cos^2(2\frac{\pi}{6})=\cos^2(\frac{\pi}{3})=(\frac{\sqrt{3}}{2})^2=\frac{3}{4}\] whereas \[\cos(4\frac{\pi}{6})=\cos(\frac{2\pi}{3})=-\frac{1}{2}\] so it is not true